What is the R code and result in Rstudio or R

1. Define a character vector in R as follows: y - c(a, bc, def) a. Report the output of the following function: length(y);ya;yb b. nchar(x) takes a character vector x as an argument and returns a vector whose elements contain the sizes of the corresponding elements of x. Report the output of nchar(y) 2. Define a numeric vector in R as follows: z=c(3,7,1,2) a. Report the output of the following function: b. which(x > 2) returns the indices of which the expression r>2 is true. Report the output of which(z <= 3) 3. To understand coercion between different data types, report the output of the following R codes a. x = c(T, F, T); c(1.2, x); 1 + x; c. y -c(3, 2); c(1.2,y); c(1.2, as.numeric(y)) 4. To understand NA (not available) in R, consider a. NA is a recognized argument in R, report the output of c(3, NA, NA);x; b. When numerical operation involves NA or missing value, the outcome will also be NA or missing. Report the output of x c(1,3, NA); mean(x) c. We can remove the missing values and then calculate the mean of the remaining elements. Report the output of x = c(1, 3, NA); mean(z, na.rm = TRUE)
Answer #1

a.

The output of the function is,

> length(y);y=="a";y=="b"
[1] 3
[1] TRUE FALSE FALSE
[1] FALSE FALSE FALSE

This shows that the length of vector y is 3.

The first element of y is "a"

None of the element of y is "b"

b.

> nchar(y)
[1] 1 2 3

The output shows that three elements of vector y are of length 1, 2 and 3.

2.

a.

The output of the function are

> (x >=2)|(x<4);(x>=2)&(x<4);!(x<3)
[1] TRUE TRUE TRUE TRUE
[1] TRUE FALSE FALSE TRUE
[1] TRUE TRUE FALSE FALSE

The output displays TRUE for the elements of x which satisfies the boolean conditions else displays FALSE.

b.

> which(x <= 3)
[1] 1 3 4

The output shows that the element at position 1, 3, 4 are less than or equal to 3.

3.

a.

> x = c(T,F,T);c(1.2,x);1+x;
[1] 1.2 1.0 0.0 1.0
[1] 2 1 2

The output shows that the code c(1.2,x) coerced the boolean variable x to numeric variable.

1 + x code coerced the real number to integer.

b.

> x = c(T,F,T);y=c("3","2");c(1.2,x,y)
[1] "1.2" "TRUE" "FALSE" "TRUE" "3" "2"

Since, R coerce boolean variable, character and numeric variable to character variable.

c.

> y=c("3","2");c(1.2,y);c(1.2,as.numeric(y))
[1] "1.2" "3" "2"
[1] 1.2 3.0 2.0

The code c(1.2,y)  coerce numeric variable to character variable.

When we convert y to numeric variable, 1.2 was not coerced to character variable.

4.

a.

> x = c(3,"NA",NA);x;
[1] "3" "NA" NA

Here "NA" is a character variable with value as NA.

NA is not applicable in R.

b.

> x = c(1,3,NA);mean(x)
[1] NA

The mean of x is NA because one of the element of x is NA.

c.

> x = c(1,3,NA);mean(x, na.rm = TRUE)
[1] 2

na.rm = TRUE parameter ignores NA value in the calculation of mean.